OpenStudy (anonymous):

The lengths of the sides of a triangle are 3, 3, 3 sqrt 2. Can the triangle be a right triangle?

OpenStudy (alexandervonhumboldt2):

to check 3^2+3^2=(3 sqrt2)^2 9+9=9*2 18=18 so yes this an be a right triangle

OpenStudy (alexandervonhumboldt2):

use pyh theorem

OpenStudy (anonymous):

Ok, hang on 1 second.

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

\[Does 3\sqrt{2} equate \to 4.2?\]

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

Does C^2 = 4.2?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

look

OpenStudy (alexandervonhumboldt2):

3^2+3^2=(3 sqrt2)^2 dont solve for sqrt2 you dont need that

OpenStudy (anonymous):

Ok, I thought the Hypotenuse had to be longer than the two legs, when added together.

OpenStudy (alexandervonhumboldt2):

(3 sqrt2)^2=3^2*2=18

OpenStudy (alexandervonhumboldt2):

18=18

OpenStudy (anonymous):

I'm slightly confused.

OpenStudy (alexandervonhumboldt2):

hmmmm pyth states that a^2+b^2=c^2 where a aand b are cathets c is the hypotenuse

OpenStudy (anonymous):

Does it become 9+9=9 sqrt2?

OpenStudy (alexandervonhumboldt2):

no 9+9=18

OpenStudy (alexandervonhumboldt2):

(3 sqrt2)^2=9*2=18 18=18

OpenStudy (alexandervonhumboldt2):

3^2+3^2=9*2 18=18

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (confluxepic):

@AlexandervonHumboldt2 You still don't understand the difference between a direct answer and a guided one. A direct answer is what you're doing (giving the whole process without the askers participation). A guided answer is when you don't say the answer. Instead you help the user with a few hints, some links, and some basic information until the user understands it. It has to have the askers participation. What you're doing is just doing all the work yourself. The asker learns nothing and takes the final result. If you want to learn more about how to give a GUIDED ANSWER look through these peoples questions and you will see. @sleepyjess @Nnesha @TheSmartOne @jigglypuff314 Look through their questions answered and learn from them so this doesn't happen again.